OpenStudy (anonymous):

Which set of equations would be used to solve this word problem? An airplane flew 5 hours with a 50 mph head wind. The return trip with a tail wind of the same speed took 3 hours. Find the speed of the plane in still air. d = 8(r + 50) d = 5(r + 50) and d = 3(r − 50) d = 5(r − 50) and d = 3(r + 50) None of these systems can be used to solve this problem.

OpenStudy (anonymous):

d? @some_someone

OpenStudy (anonymous):

d = 5(r + 50) and d = 3(r - 50)